OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (aakashsudhakar):

These are all in point-slope form. Remember that point-slope form is represented by y - y1 = m(x - x1), where (x1, y1) are any given coordinate on the function. You're given that m = -6 and that (x1, y1) = (-2, -3). Simply plug them into the point-slope formula to find the appropriate equation. y - y1 = m(x - x1) y + 3 = -6(x + 2) In this case, the correct answer would be the first option.